# Jinhong Li1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

### v2

@@ 5.2 Microbenchmarks
(a) Read throughput (b) Median read latency (c) 95p read latency
Figure 7: Exp#2 (Normal and degraded read performance). We consider normal reads (NR), degraded reads under static mapping in Log-RAID (DR-Log-RAID), and degraded reads under the group-based data layout in ZapRAID (DR-ZapRAID).
Exp#3 (Impact of stripe group size).
We study the impact of the stripe group size $G$ on the write and read performance in ZapRAID. We vary $G$ from 4 to 4,096. For both writes and reads, we vary the chunk size and set the request size to be the same as the chunk size.
Figure 8 shows the write throughput and the median degraded read latency of ZapRAID versus $G$. The write throughput of ZapRAID increases with $G$ for $G\leq 256$ and remains stable for $G\geq 256$ under 4-KiB and 8-KiB requests, while it remains almost unchanged under 16-KiB requests. The write throughput reaches 1,574.2 MiB/s and 2,961.2 MiB/s for $G=256$ under the chunk size of 4 KiB and 8 KiB (1.43$\times$ and 1.59$\times$ the write throughput for $G=4$), respectively. Thus, ZapRAID can better exploit the intra-zone parallelism through Zone Append as $G$ increases, but the parallelism saturates when $G$ becomes sufficiently large. For 16-KiB chunks, the intra-zone parallelism brought by Zone Append is limited and already saturates when $G$ is small. For the degraded read latency, when $G$ is small, the query overhead of the compact stripe table is negligible (around 1 $\mu$s based on our measurement). However, as $G$ continues to increase, the query overhead becomes non-negligible. For example, the degraded read latency for $G=$ 4,096 increases by 13.2-24.7% compared with that for $G=256$. In addition to I/O performance, a large $G$ also increases the memory usage of the compact stripe table (§3.2). Thus, we choose $G=256$ as our default setting.
@@
(a) Write throughput (b) Median degraded read latency
Figure 8: Exp#3 (Impact of stripe group size).
We also examine the degraded read latency of ZoneAppend-Only (not shown in the figure). We find that its median degraded latency is 1.84 ms (i.e., 21.6$\times$ compared with ZapRAID for $G=256$). Thus, ZoneAppend-Only not only incurs high memory usage (§3.2), but also has poor degraded read performance due to its high query overhead to the compact stripe table.
Exp#4 (Impact of RAID schemes).
ZapRAID supports various RAID schemes. We configure RAID-0, RAID-01, RAID-4, RAID-5, and RAID-6 on the four ZNS SSDs (e.g., $k=2$ and $m=2$ in RAID-6). Here, we compare ZapRAID and ZoneWrite-Only. We fix the request size to be the same as the chunk size and vary the chunk size with a queue depth of 64.
Figure 9 depicts the results. ZapRAID increases the write throughput of ZoneWrite-Only by 71.5-72.1%, 75.8-76.4%, and 5.3-5.7% for 4-KiB, 8-KiB, and 16-KiB chunks, respectively. The difference in the write throughput among the RAID schemes is due to the different number of data chunks in each stripe (i.e., four, two, three, three, and two for RAID-0, RAID-01, RAID-4, RAID-5, and RAID-6, respectively). The results show that ZapRAID consistently increases the write throughput of ZoneWrite-Only under different RAID schemes.
@@
(a) 4-KiB chunks (b) 8-KiB chunks (c) 16-KiB chunks
Figure 9: Exp#4 (Impact of RAID schemes).
Exp#5 (Recovery performance).
We evaluate the recovery performance of ZapRAID in two aspects: recovery from a system crash (§3.4) and recovery from a full-drive failure (§3.5). We sequentially write data to ZapRAID configured with a fixed size of logical space (varying from 100 GiB to 1,000 GiB), mimic each type of failure, and repeat the recovery procedure for five runs.
We first evaluate the crash recovery time of ZapRAID. We crash ZapRAID and rebuild it from the currently stored data in the drives. We report the average recovery time (i.e., the time spent on performing all recovery operations in §3.4). Figure 10(a) depicts the results for different chunk sizes. For the 100-GiB storage space, ZapRAID takes 2.38-2.60 s to recover the system state. The recovery time increases linearly with the logical space size. The reason is that for larger logical space, ZapRAID needs to recover the system state from more sealed segments by reading their footer regions. For example, for the 1,000-GiB storage space, the recovery time increases to 15.92-16.22 s. Nevertheless, the recovery overhead is limited. For example, for 4-KiB chunks, we only need to pay an additional $(16.22-2.60)/(10-1)=1.51$ s for recovering the system state from each 100 GiB of additional logical space. Also, the crash recovery time remains almost the same for different chunk sizes, since most of the time is spent on reading the footer regions with a large read size, independent of the chunk size.
We next evaluate the full-drive recovery time of ZapRAID. We erase all data in one drive and recover the lost data in the same drive. We report the average recovery time as the total time spent on reading available data from existing drives, reconstructing the stripes, and writing the recovered data to the new drive. Figure 10(b) shows the results for different chunk sizes. The full-drive recovery time is proportional to the logical space size. For example, for 4-KiB chunks, ZapRAID takes 81.9 s and 813.2 s to recover the lost data for the 100-GiB and 1,000-GiB storage space, respectively. On average, for each 100-GiB storage space, ZapRAID spends $813.2/10=81.3$ s to recover the lost data. Note that the recovery time is smaller for larger chunk sizes since ZapRAID can write recovered data with a larger write size. For example, for 8-KiB and 16-KiB chunks, the recovery time is reduced by 18.1-22.2% and 22.1-23.8% compared with 4-KiB chunks, respectively.
@@
Figure 10: Exp#5 (Recovery performance).
Exp#6 (Scalability).
We examine the scalability of ZapRAID by evaluating its write performance under different values of queue depth and showing how it saturates the system parallelism with a high queue depth. We consider both real and FEMU-emulated [53] ZNS SSDs.
@@
Figure 11(a) depicts the results. ZapRAID increases the write throughput when the queue depth increases. It saturates the system parallelism when the queue depth is 16, in which the throughput gain is 3.52$\times$, 3.57$\times$, and 2.08$\times$ compared with when the queue depth is four for 4-KiB, 8-KiB, and 16-KiB chunks, respectively. Note that we set the queue depth for all microbenchmark experiments as 64 to further increase the throughput for multiple open segments (Exp#7).
Our current testbed has limited parallelism as it only has four real ZNS SSDs. To examine the scalability of ZapRAID with a higher degree of parallelism, we have also built an emulated ZNS SSD array with FEMU [53]. We set up a (6+1)-RAID-5 array on seven FEMU SSDs, using 4$\times$128-GiB Intel Optane Persistent Memory Modules [2] in Memory Mode as the emulated persistent storage backend. Each FEMU SSD is configured with four channels with four chips each. Each chip has 64 flash blocks with 4,096 4-KiB flash pages each. Each zone spans 16 chips, meaning that I/Os can be issued to up to 16 chips in parallel. We configure the page write latency as 140 $\mu$s [54] and do not emulate write buffers in SSDs that can reduce write latency. The page write latency translates to the write throughput of 27.9 MiB/s per chip, or equivalently, 446.4 MiB/s per zone if the parallelism is fully exploited. Since the FEMU SSD does not support flash-level page metadata, we store the page metadata in an in-memory array. We configure seven FEMU SSDs in a FEMU virtual machine, installed with Ubuntu 22.04, 32-GiB memory, and 12 vCPUs. We export the ZapRAID prototype as a block device with 96 GiB of logical space. We use FIO to issue the random writes of 64 GiB of data with a block size of 4 KiB to ZapRAID in the virtual machine for various chunk sizes and queue depths.
Figure 11(b) shows the results. When the queue depth is four, the write throughput for 4-KiB, 8-KiB, and 16-KiB chunks are 58.3 MiB/s, 114.0 MiB/s, and 224.2 MiB/s, respectively. As the queue depth increases, the write throughput of ZapRAID increases and achieves 1,745.5 MiB/s, 2,384.6 MiB/s, and 2,554.2 MiB/s (i.e., 29.9$\times$, 20.9$\times$, and 11.4$\times$ compared with the queue depth of four), respectively when the queue depth is 128.
@@
Figure 11: Exp#6 (Scalability). The results in figures (a) and (b) are based on ZapRAID, and the results in figure (c) are based on the conventional RAID array.
Exp#7 (Multiple open segments).
We evaluate the write performance ZapRAID on multiple open segments based on hybrid data management (§3.3) by varying the numbers of small-chunk and large-chunk segments (i.e., $N_{s}$ and $N_{l}$, respectively). We fix $C_{s}=$ 8 KiB, $C_{l}=$ 16 KiB, and the number of open segments as four. We consider different combinations of $N_{s}$ and $N_{l}$, such that $N_{s}+N_{l}=4$. Note that if $N_{s}=0$ (resp. $N_{l}=0$), all writes are issued to large-chunk (resp. small-chunk) segments.
We consider four workloads. The first three workloads issue 4-KiB, 8-KiB, and 16-KiB writes via FIO, respectively. Note that these workloads do not use all open segments since they write 4-KiB or 8-KiB requests to small-chunk segments only and 16-KiB requests to large-chunk segments only. The fourth workload consists of a mixture of 75% 4-KiB writes and 25% 16-KiB writes by setting the bssplit parameter in FIO, based on the observation that 25% of write requests in real-world cloud block storage workloads are at least 16 KiB and most of the write requests are smaller than 4 KiB [55]. We set the total write size of each workload as 64 GiB.
Figures 12(a)-12(d) show the throughput results. ZoneWrite-Only and ZoneAppend-Only cannot always achieve the highest throughput in all cases. For example, ZoneAppend-Only has 65.7% higher throughput than ZoneWrite-Only for 4-KiB writes and $(N_{s},N_{l})=(1,3)$ due to better intra-zone parallelism under small writes (Figure 12(a)), but ZoneWrite-Only has 27.2% higher throughput than ZoneAppend-Only for 16-KiB writes and $(N_{s},N_{l})=(1,3)$ (Figure 12(c)). On the other hand, ZapRAID achieves the highest throughput, or similar throughput to the best scheme, in all cases by carefully combining Zone Append and Zone Write.
Note that while ZoneAppend-Only can issue 16-KiB writes to multiple large-chunk segments, it cannot achieve higher throughput as $N_{l}$ increases, which may be inconsistent with the findings in Figure 22.1). One possible reason is that the computational overhead of the firmware for Zone Append becomes heavier and has higher fluctuations when we issue writes via Zone Append to multiple drives, and the performance is bottlenecked by the current slowest drive. Note that Zone Write also has a similar problem, but is less severe than Zone Append. Such performance fluctuations are also observed in prior work [76, 38].
@@
(a) 4-KiB writes, throughput (b) 8-KiB writes, throughput
@@
(c) 16-KiB writes, throughput (d) Hybrid 4-KiB and 16-KiB writes, throughput
@@
(e) 4-KiB writes, 95p latency (f) 8-KiB writes, 95p latency
@@
(g) 16-KiB writes, 95p latency (h) Hybrid 4-KiB and 16-KiB writes, 95p latency
@@
Figure 12: Exp#7 (Multiple open segments). We show the throughput (figures (a)-(d)) and 95th-percentile latencies (figures (e)-(h)) under four open segments.
@@
Figures 12(e)-12(h) show the 95th-percentile latency results. Overall, ZapRAID maintains low 95th-percentile latencies in all cases. For 4-KiB writes (Figure 12(e)) and hybrid 4-KiB and 16-KiB writes (Figure 12(h)), all three schemes have similar latencies except for $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only has a significantly high latency since the 4-KiB writes are issued to the only small-chunk segment and ZoneWrite-Only has limited intra-zone parallelism.
@@
For 8-KiB writes (Figure 12(f)), ZapRAID and ZoneWrite-Only have similar 95th-percentile latencies, while ZapRAID has 15.9-32.7% lower 95th-percentile latencies than ZoneAppend-Only for $(N_{s},N_{l})=$ (4,0), (3,1), and (2,2) by better exploiting inter-zone parallelism through both Zone Append and Zone Write. For $(N_{s},N_{l})=(1,3)$, ZapRAID has a slightly higher 95th-percentile latency than ZoneAppend-Only (by 9.4%) due to the overhead of group-based data layout.
@@
For 16-KiB writes (Figure 12(g)), ZapRAID and ZoneWrite-Only have consistently lower 95th-percentile latencies than ZoneAppend-Only. We observe that the 95th-percentile latency of ZoneAppend-Only grows sharply by 173.3% from $(N_{s},N_{l})=(4,0)$ to $(N_{s},N_{l})=(3,1)$ since the 16-KiB writes are issued to the only open large-chunk segment for $N_{l}=1$ while they are issued to four open small-chunk segments for $N_{l}=0$. The 95th-percentile latency of issuing 16-KiB writes to a single large-chunk segment is much higher (see also Figure 6(c) in Exp#1). As $N_{l}$ further increases from one to four, the 95th-percentile latency of ZoneAppend-Only drops by 58.3% as there are more open large-chunk segments and a higher degree of inter-zone parallelism. We also observe that the 95th-percentile latencies of ZapRAID and ZoneWrite-Only increase by 121.1% and 115.3%, respectively from $(N_{s},N_{l})=(3,1)$ to $(N_{s},N_{l})=(2,2)$. The reason is that the Zone Write command on one open segment has low on-SSD computational overhead as it only handles one outstanding write request. The overhead grows when it issues multiple outstanding write requests to different open segments. Similar to ZoneAppend-Only, their 95th-percentile latencies drop as $N_{l}$ increases from two to four due to a higher degree of inter-zone parallelism.
@@
(c) $N_{l}=2$ and varying $N_{s}$, 95p latency (d) $N_{s}=2$ and varying $N_{l}$, 95p latency
Figure 13: Exp#7 (Multiple open segments). We compare the throughput and 95th-percentile latencies of RAIZN-SPDK and ZapRAID under hybrid 75% 4-KiB and 25% 16-KiB writes and different combinations of $(N_{s},N_{l})$.
We also compare ZapRAID with RAIZN-SPDK under multiple open segments. We focus on the workloads with hybrid 75% 4-KiB and 25% 16-KiB writes, and examine different combinations of $(N_{s},N_{l})$. We consider two settings: (i) we fix $N_{l}=2$ and vary $N_{s}$ from 0 to 6, and (ii) we fix $N_{s}=2$ and vary $N_{l}$ from 0 to 6.
Figure 13(a) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{l}=2$ and vary $N_{s}$. For RAIZN-SPDK, when $(N_{s},N_{l})$ changes from $(0,2)$ to $(1,2)$, its throughput drops from 440.4 MiB/s to 268.6 MiB/s. The reason is that RAIZN-SPDK forwards all 4-KiB writes to two large-chunk segments for $(N_{s},N_{l})=(0,2)$, but for $(N_{s},N_{l})=(1,2)$, it forwards all 4-KiB writes to only one small-chunk segment and fails to leverage inter-zone parallelism for these small writes. When $(N_{s},N_{l})$ changes from $(1,2)$ to $(6,2)$, the throughput of RAIZN-SPDK increases to 1,077.6 MiB/s with the increasing inter-zone parallelism across small-chunk open segments. For ZapRAID, with the help of intra-zone parallelism for small writes, it maintains similar throughput for $(N_{s},N_{l})=(0,2)$ and $(N_{s},N_{l})=(1,2)$, even though the latter has only one small-chunk open segment.
Figure 13(b) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{s}=2$ and vary $N_{l}$. The throughput of RAIZN-SPDK is 444.6 MiB/s for $(N_{s},N_{l})=(2,0)$, but only increases to 488.4-498.0 MiB/s when $(N_{s},N_{l})$ varies from $(2,1)$ to $(2,6)$. The reason is that most of the write requests (75% in total) are 4-KiB writes, and the performance is bottlenecked by the limited number of small-chunk open segments.
Figures 13(c) and 13(d) show the 95th-percentile latency results of RAIZN-SPDK and ZapRAID, and the observations are consistent with those in the throughput results. Overall, when there are multiple open segments, ZapRAID still achieves significantly higher throughput and lower latencies than RAIZN-SPDK, which incurs large overhead of partial parity updates.
@@
Table 1: Exp#7 (Multiple open segments). Breakdown of latencies (in microseconds) of ZapRAID and RAIZN-SPDK for Figure 13.
@@
(a) 4-KiB chunks (b) 16-KiB chunks (c) Multiple open segments
Figure 14: Exp#8 (Overhead of garbage collection).
Exp#8 (Overhead of garbage collection in ZapRAID).
@@
We evaluate the impact of garbage collection on ZapRAID with different reserved space sizes. We issue writes to 200 GiB of data, with a total of 1 TiB of write traffic. We vary the physical space capacity of ZapRAID from 240 GiB to 400 GiB, meaning that the reserved space for garbage collection varies between 20% and 100%. We consider four write workloads, including 4-KiB (uniform) random writes, 64-KiB (uniform) random writes, 64-KiB skewed writes, and 64-KiB sequential writes. We consider 64-KiB writes to examine the overhead of garbage collection caused by much larger writes. For 64-KiB skewed writes, we consider a Zipfian distribution with a skewness factor of 0.99. We consider three settings: (i) a single open segment with 4-KiB chunks, (ii) a single open segment with 16-KiB chunks, and (iii) four open segments with $(N_{s},N_{l})=(2,2)$, $C_{s}=$ 8 KiB, and $C_{l}=$ 16 KiB.
@@
Figure 14 shows the results. As expected, the write throughput increases (i.e., the garbage collection overhead is less) with more reserved space and with more skewed or sequential workloads. Since real-world workloads are often skewed in practice (e.g., block storage [74] and key-value [18, 24] workloads), we expect that the garbage collection overhead is acceptable in ZapRAID. Note that the write throughput under multiple open segments (Figure 14(c)) is higher than under single open segments (Figures 14(a) and 14(b)), meaning that the garbage collection overhead can be reduced via inter-zone parallelism.
Exp#9 (Overhead of L2P table offloading).

### v3

@@ 5.4 Microbenchmarks (RQ3)
(a) Read throughput (b) Median read latency (c) p95 read latency
Figure 11: Exp#6 (Normal and degraded read performance). We consider normal reads (NR), degraded reads under static mapping in Log-RAID (DR-Log-RAID), and degraded reads under the group-based data layout in ZapRAID (DR-ZapRAID).
Exp#7 (Impact of stripe group size). We study the impact of the stripe group size $G$ on the write and read performance in ZapRAID. We vary $G$ from 4 to 4,096. For both writes and reads, we vary the chunk size and set the request size to be the same as the chunk size.
[…]
@@
(a) Write throughput for ZapRAID (b) Median degraded read latencies for different schemes
Figure 12: Exp#7 (Impact of stripe group size).
[…]
Exp#8 (Impact of RAID schemes). ZapRAID supports various RAID schemes. We configure RAID-0, RAID-01, RAID-4, RAID-5, and RAID-6 on the four ZNS SSDs (e.g., $k=2$ and $m=2$ in RAID-6). Here, we compare ZapRAID and ZoneWrite-Only. We fix the request size to be the same as the chunk size and vary the chunk size with a queue depth of 64.
Figure 13 depicts the results. ZapRAID increases the write throughput of ZoneWrite-Only by 71.5-72.1%, 75.8-76.4%, and 5.3-5.7% for 4 KiB, 8 KiB, and 16 KiB chunks, respectively. The difference in the write throughput among the RAID schemes is due to the different number of data chunks in each stripe (i.e., four, two, three, three, and two for RAID-0, RAID-01, RAID-4, RAID-5, and RAID-6, respectively). The results show that ZapRAID consistently increases the write throughput of ZoneWrite-Only under different RAID schemes.
@@
(a) 4 KiB chunks (b) 8 KiB chunks (c) 16 KiB chunks
Figure 13: Exp#8 (Impact of RAID schemes).
Exp#9 (Recovery performance). We evaluate the recovery performance of ZapRAID in two aspects: recovery from a system crash (§3.4) and recovery from a full-drive failure (§3.5). We sequentially write data to ZapRAID configured with a fixed size of logical space (varying from 100 GiB to 1,000 GiB), mimic each type of failure, and repeat the recovery procedure for five runs.
We first evaluate the crash recovery time of ZapRAID. We crash ZapRAID and rebuild it from the currently stored data in the drives. We report the average recovery time (i.e., the time spent on performing all recovery operations in §3.4). Figure 14(a) depicts the results for different chunk sizes. For the 100 GiB storage space, ZapRAID takes 2.38-2.60 s to recover the system state. The recovery time increases linearly with the logical space size. The reason is that for larger logical space, ZapRAID needs to recover the system state from more sealed segments by reading their footer regions. For example, for the 1,000 GiB storage space, the recovery time increases to 15.92-16.22 s. Nevertheless, the recovery overhead is limited. For example, for 4 KiB chunks, we only need to pay an additional (16.222.60)/(101)=1.51(16.22-2.60)/(10-1)=1.51 s for recovering the system state from each 100 GiB of additional logical space. Also, the crash recovery time remains almost the same for different chunk sizes, since most of the time is spent on reading the footer regions with a large read size, independent of the chunk size.
We next evaluate the full-drive recovery time of ZapRAID. We erase all data in one drive and recover the lost data in the same drive. We report the average recovery time as the total time spent on reading available data from existing drives, reconstructing the stripes, and writing the recovered data to the new drive. Figure 14(b) shows the results for different chunk sizes. The full-drive recovery time is proportional to the logical space size. For example, for 4 KiB chunks, ZapRAID takes 81.9 s and 813.2 s to recover the lost data for the 100 GiB and 1,000 GiB storage space, respectively. On average, for each 100 GiB storage space, ZapRAID spends $813.2/10=81.3$ s to recover the lost data. Note that the recovery time is smaller for larger chunk sizes since ZapRAID can write recovered data with a larger write size. For example, for 8 KiB and 16 KiB chunks, the recovery time is reduced by 18.1-22.2% and 22.1-23.8% compared with 4 KiB chunks, respectively.
@@
Figure 14: Exp#9 (Recovery performance).
Exp#10 (Scalability). We examine the scalability of ZapRAID by evaluating its write performance under different values of queue depth and showing how it saturates the system parallelism with a high queue depth. We consider both real and FEMU-emulated [39] ZNS SSDs.
@@
[…]
@@
Figure 15: Exp#10 (Scalability). The results in figures (a) and (b) are based on ZapRAID, and the results in figure (c) are based on the conventional RAID array.
Exp#11 (Multiple open segments). We evaluate the write performance ZapRAID on multiple open segments based on hybrid data management (§3.3) by varying the numbers of small-chunk and large-chunk segments (i.e., $N_{s}$ and $N_{l}$, respectively). As introduced in §5.2, we fix $C_{s}=$ 8 KiB, $C_{l}=$ 16 KiB, and the number of open segments as four. We consider different combinations of $N_{s}$ and $N_{l}$, such that $N_{s}+N_{l}=4$. Note that if $N_{s}=0$ (resp. $N_{l}=0$), all writes are issued to large-chunk (resp. small-chunk) segments.
We consider four workloads. The first three workloads issue 4 KiB, 8 KiB, and 16 KiB writes via FIO, respectively. Note that these workloads do not use all open segments since they write 4 KiB or 8 KiB requests to small-chunk segments only and 16 KiB requests to large-chunk segments only. The fourth workload consists of a mixture of 75% 4 KiB writes and 25% 16 KiB writes by setting the bssplit parameter in FIO, based on the observation that 25% of write requests in real-world cloud block storage workloads are at least 16 KiB and most of the write requests are smaller than 4 KiB [41]. We set the total write size of each workload as 64 GiB.
Figures 16(a)-16(d) show the throughput results. ZoneWrite-Only and ZoneAppend-Only cannot always achieve the highest throughput in all cases. For example, ZoneAppend-Only has 65.7% higher throughput than ZoneWrite-Only for 4 KiB writes and $(N_{s},N_{l})=(1,3)$ due to better intra-zone parallelism under small writes (Figure 16(a)), but ZoneWrite-Only has 27.2% higher throughput than ZoneAppend-Only for 16 KiB writes and $(N_{s},N_{l})=(1,3)$ (Figure 16(c)). On the other hand, ZapRAID achieves the highest throughput, or similar throughput to the best scheme, in all cases by carefully combining Zone Append and Zone Write.
Note that while ZoneAppend-Only can issue 16 KiB writes to multiple large-chunk segments, it cannot achieve higher throughput as $N_{l}$ increases, which may be inconsistent with the findings in Figure 22.1). One possible reason is that the computational overhead of the firmware for Zone Append becomes heavier and has higher fluctuations when we issue writes via Zone Append to multiple drives, and the performance is bottlenecked by the current slowest drive. Note that Zone Write also has a similar problem, but is less severe than Zone Append. Such performance fluctuations are also observed in prior work [75, 24].
@@
(a) 4 KiB writes, throughput (b) 8 KiB writes, throughput
@@
(c) 16 KiB writes, throughput (d) Hybrid 4 KiB and 16 KiB writes, throughput
@@
(e) 4 KiB writes, p95 latency (f) 8 KiB writes, p95 latency
@@
(g) 16 KiB writes, p95 latency (h) Hybrid 4 KiB and 16 KiB writes, p95 latency
Refer to caption Refer to caption
(i) 4 KiB writes, memory size (j) 8 KiB writes, memory size
Refer to caption Refer to caption
(k) 16 KiB writes, memory size (l) Hybrid 4 KiB and 16 KiB writes, memory size
@@
Figure 16: Exp#11 (Multiple open segments). We show the throughput (figures (a)-(d)), 95th-percentile latencies (figures (e)-(h)), and stripe management memory usage per GiB of written data (figures (i)-(l)) under four open segments.
@@
Figures 16(e)-16(h) show the 95th-percentile latency results. Overall, ZapRAID maintains low 95th-percentile latencies in all cases. For 4 KiB writes (Figure 16(e)) and hybrid 4 KiB and 16 KiB writes (Figure 16(h)), all three schemes have similar latencies except for $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only has a significantly high latency since the 4 KiB writes are issued to the only small-chunk segment and ZoneWrite-Only has limited intra-zone parallelism.
@@
For 8 KiB writes (Figure 16(f)), ZapRAID and ZoneWrite-Only have similar 95th-percentile latencies, while ZapRAID has 15.9-32.7% lower 95th-percentile latencies than ZoneAppend-Only for $(N_{s},N_{l})=$ (4,0), (3,1), and (2,2) by better exploiting inter-zone parallelism through both Zone Append and Zone Write. For $(N_{s},N_{l})=(1,3)$, ZapRAID has a slightly higher 95th-percentile latency than ZoneAppend-Only (by 9.4%) due to the overhead of group-based data layout.
@@
For 16 KiB writes (Figure 16(g)), ZapRAID and ZoneWrite-Only have consistently lower 95th-percentile latencies than ZoneAppend-Only. We observe that the 95th-percentile latency of ZoneAppend-Only grows sharply by 173.3% from $(N_{s},N_{l})=(4,0)$ to $(N_{s},N_{l})=(3,1)$ since the 16 KiB writes are issued to the only open large-chunk segment for $N_{l}=1$ while they are issued to four open small-chunk segments for $N_{l}=0$. The 95th-percentile latency of issuing 16 KiB writes to a single large-chunk segment is much higher (see also Figure 9(c) in Exp#5). As $N_{l}$ further increases from one to four, the 95th-percentile latency of ZoneAppend-Only drops by 58.3% as there are more open large-chunk segments and a higher degree of inter-zone parallelism. We also observe that the 95th-percentile latencies of ZapRAID and ZoneWrite-Only increase by 121.1% and 115.3%, respectively from $(N_{s},N_{l})=(3,1)$ to $(N_{s},N_{l})=(2,2)$. The reason is that the Zone Write command on one open segment has low on-SSD computational overhead as it only handles one outstanding write request. The overhead grows when it issues multiple outstanding write requests to different open segments. Similar to ZoneAppend-Only, their 95th-percentile latencies drop as $N_{l}$ increases from two to four due to a higher degree of inter-zone parallelism.
Figures 16(i)-16(l) show the stripe management memory size per GiB of written data. We focus on ZoneAppend-Only and ZapRAID as in Exp#5. ZapRAID maintains significantly lower memory usage with its compact stripe table design than ZoneAppend-Only in all cases. For example, for 4 KiB writes (Figure 16(i)), ZapRAID achieves 68.9-98.5% lower memory usage than ZoneAppend-Only for $(N_{s},N_{l})=(4,0)$, $(3,1)$, $(2,2)$, and $(1,3)$. Also, ZapRAID incurs no memory usage for the compact stripe table for the cases of (i) $(N_{s},N_{l})=(0,4)$ and (ii) $N_{l}>0$ and there are only large writes (i.e., 16 KiB writes), since it only issues Zone Write to large-chunk segments. Thus, ZapRAID keeps low or no memory usage for stripe management, while ZoneAppend-Only always incurs significant memory usage for tracking the mappings between stripes and chunk locations.
@@
(c) $N_{l}=2$ and varying $N_{s}$, p95 latency (d) $N_{s}=2$ and varying $N_{l}$, p95 latency
Figure 17: Exp#11 (Multiple open segments). We compare the throughput and 95th-percentile latencies of RAIZN-SPDK and ZapRAID under hybrid 75% 4 KiB and 25% 16 KiB writes and different combinations of $(N_{s},N_{l})$.
We also compare ZapRAID with RAIZN-SPDK under multiple open segments. We focus on the workloads with hybrid 75% 4 KiB and 25% 16 KiB writes, and examine different combinations of $(N_{s},N_{l})$. We consider two settings: (i) we fix $N_{l}=2$ and vary $N_{s}$ from 0 to 6, and (ii) we fix $N_{s}=2$ and vary $N_{l}$ from 0 to 6.
Figure 17(a) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{l}=2$ and vary $N_{s}$. For RAIZN-SPDK, when $(N_{s},N_{l})$ changes from $(0,2)$ to $(1,2)$, its throughput drops from 440.4 MiB/s to 268.6 MiB/s. The reason is that RAIZN-SPDK forwards all 4 KiB writes to two large-chunk segments for $(N_{s},N_{l})=(0,2)$, but for $(N_{s},N_{l})=(1,2)$, it forwards all 4 KiB writes to only one small-chunk segment and fails to leverage inter-zone parallelism for these small writes. When $(N_{s},N_{l})$ changes from $(1,2)$ to $(6,2)$, the throughput of RAIZN-SPDK increases to 1,077.6 MiB/s with the increasing inter-zone parallelism across small-chunk open segments. For ZapRAID, with the help of intra-zone parallelism for small writes, it maintains similar throughput for $(N_{s},N_{l})=(0,2)$ and $(N_{s},N_{l})=(1,2)$, even though the latter has only one small-chunk open segment.
Figure 17(b) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{s}=2$ and vary $N_{l}$. The throughput of RAIZN-SPDK is 444.6 MiB/s for $(N_{s},N_{l})=(2,0)$, but only increases to 488.4-498.0 MiB/s when $(N_{s},N_{l})$ varies from $(2,1)$ to $(2,6)$. The reason is that most of the write requests (75% in total) are 4 KiB writes, and the performance is bottlenecked by the limited number of small-chunk open segments.
Figures 17(c) and 17(d) show the 95th-percentile latency results of RAIZN-SPDK and ZapRAID, and the observations are consistent with those in the throughput results. Overall, when there are multiple open segments, ZapRAID still achieves significantly higher throughput and lower latencies than RAIZN-SPDK, which incurs large overhead of partial parity updates.
@@
Table 4: Exp#11 (Multiple open segments). Breakdown of latencies (in microseconds) of ZapRAID and RAIZN-SPDK for Figure 17.
@@
(a) 4 KiB chunks (b) 16 KiB chunks (c) Multiple open segments
Figure 18: Exp#12 (Overhead of garbage collection).
@@
Exp#12 (Overhead of garbage collection in ZapRAID). We evaluate the impact of garbage collection on ZapRAID with different reserved space sizes. We issue writes to 200 GiB of data, with a total of 1 TiB of write traffic. We vary the physical space capacity of ZapRAID from 240 GiB to 400 GiB, meaning that the reserved space for garbage collection varies between 20% and 100%. We consider four write workloads, including 4 KiB (uniform) random writes, 64 KiB (uniform) random writes, 64 KiB skewed writes, and 64 KiB sequential writes. We consider 64 KiB writes to examine the overhead of garbage collection caused by much larger writes. For 64 KiB skewed writes, we consider a Zipfian distribution with a skewness factor of 0.99. We consider three settings: (i) a single open segment with 4 KiB chunks, (ii) a single open segment with 16 KiB chunks, and (iii) four open segments with $(N_{s},N_{l})=(2,2)$, $C_{s}=$ 8 KiB, and $C_{l}=$ 16 KiB.
@@
Figure 18 shows the results. As expected, the write throughput increases (i.e., the garbage collection overhead is less) with more reserved space and with more skewed or sequential workloads. Since real-world workloads are often skewed in practice (e.g., block storage [72] and key-value [2, 8] workloads), we expect that the garbage collection overhead is acceptable in ZapRAID. Note that the write throughput under multiple open segments (Figure 18(c)) is higher than under single open segments (Figures 18(a) and 18(b)), meaning that the garbage collection overhead can be reduced via inter-zone parallelism.